Seminar on mirror symmetry and representation theory

(upgraded from “Seminar on derived categories in Algebraic Geometry”)

Tuesdays, 5.00-6.30, MIT, room 2-132.

Tuesday, February 28

Homological mirror symmetry for ℂℙ^{2} (continuation).

by Denis Auroux (MIT).

Denis will continue his explanation of Homological Mirror Symmetry for Fano toric varieties on the example of ℂℙ^{2}.

seminar contacts: Roman Bezrukavnikov (bezrukav@math.mit.edu); Rina Anno (rina@math.harvard.edu)